current age of the children is supposed as h
In ten years -> h+10, a child will be three years older than twice her current age -> 2h+3.
h+10 = 2h+3
what is the current age of the child? h = ?

Answer by amalialaila(3)

h+10 = 2h+3
h-2h = 3-10
-h = -7 ---> multiply -1
h = 7
the child's current age is 7.